# Long-time behavior for discretization schemes of Fokker–Planck equations via couplings

### v1

@@ Appendix A Convergence of the Markov chain to the SDE
We prove Theorem 3. The proof relies on the diffusion approximation of [17, Theorem 7.4.1], and it is rather standard. We present it for completeness. For $n\in{\mathbb{N}}$, let $h=2^{-n}\xi$, where $\xi>0$ is such that $2K/h\in{\mathbb{N}}$ holds. We write $(Y_{t}^{n})$ instead of $(Y_{t}^{h})$, ${\mathcal{L}}_{n}$ instead of ${\mathcal{L}}_{h}$, ${\mathcal{K}}_{n}$ instead of ${\mathcal{K}}_{h}$, and $V^{n}$ instead of $V^{h}$. For each $f\in\mathcal{A}$ with $\mathcal{A}$ defined in (4), we introduce the process
@@
×$\phantom{xx}\times\big{(}f(Y_{s}^{n}\pm n^{-1}e_{j})-f(Y_{s}^{n})\big{)}{\mathrm{d}}s\bigg{|},$
@@
×$\phantom{xx}\times\big{(}f(Y_{s}^{n}\pm n^{-1}e_{j})-f(Y_{s}^{n})\big{)}{\mathrm{d}}s\bigg{|},$
@@
×$\phantom{xx}\times\partial_{j}^{\pm}f(Y_{s}^{n})\partial_{j}^{\pm}g(Y_{s}^{n}){\mathrm{d}}s\bigg{|},$
@@
Summarizing the previous results and taking into account [17, Theorem 7.4.1], we conclude that the sequence of processes $(M_{f}^{n})_{n\in{\mathbb{N}}}$ and also $(B_{f}^{n})_{n\in{\mathbb{N}}}$ are relatively compact in the space of càdlàg functions $f:{\mathbb{R}}_{+}^{d}\to D$, i.e. of right-continuous functions for which the left limit exists. Hence, $(f(Y_{t}^{n}))_{t\geq 0}$ is relatively compact for all $f\in\mathcal{A}$. Then, for given $f\in\mathcal{A}$, there exists a subsequence $(n_{k})_{k\in{\mathbb{N}}}$ such that $f(Y_{t}^{n_{k}})_{t\geq 0}$ converges to some limit $(X_{t}^{f})_{t\geq 0}$.
@@
Clearly, $(M_{f}^{\bar{n}_{k}})_{k}$, $(B_{f}^{\bar{n}_{k}})_{k}$, and $f(Y_{t}^{\bar{n}_{k}})_{k}$ are relatively compact for $f\in\mathcal{A}$. As in [17, Theorem 7.4.1], $(M_{f}^{n_{k}})_{k}$ is uniformly integrable for $f\in\mathcal{A}$. Consequently,

### v2

@@ Appendix A Convergence of the Markov chain to the SDE
We prove Theorem 3. The proof relies on the diffusion approximation of [18, Theorem 7.4.1], and it is rather standard. We present it for completeness. For $n\in{\mathbb{N}}$, let $h=2^{-n}\xi$, where $\xi>0$ is such that $2K/h\in{\mathbb{N}}$ holds. We write $(Y_{t}^{n})$ instead of $(Y_{t}^{h})$, ${\mathcal{L}}_{n}$ instead of ${\mathcal{L}}_{h}$, ${\mathcal{K}}_{n}$ instead of ${\mathcal{K}}_{h}$, and $V^{n}$ instead of $V^{h}$. For each $f\in\mathcal{A}$ with $\mathcal{A}$ defined in (4), we introduce the process
@@
$\phantom{xx}\times\big(f(Y_{s}^{n}\pm n^{-1}e_{j})-f(Y_{s}^{n})\big){\mathrm{d}}s\bigg|,$
@@
$\phantom{xx}\times\big(f(Y_{s}^{n}\pm n^{-1}e_{j})-f(Y_{s}^{n})\big){\mathrm{d}}s\bigg|,$
@@
$\phantom{xx}\times\partial_{j}^{\pm}f(Y_{s}^{n})\partial_{j}^{\pm}g(Y_{s}^{n}){\mathrm{d}}s\bigg|,$
@@
Summarizing the previous results and taking into account [18, Theorem 7.4.1], we conclude that the sequence of processes $(M_{f}^{n})_{n\in{\mathbb{N}}}$ and also $(B_{f}^{n})_{n\in{\mathbb{N}}}$ are relatively compact in the space of càdlàg functions $f:{\mathbb{R}}_{+}^{d}\to D$, i.e. of right-continuous functions for which the left limit exists. Hence, $(f(Y_{t}^{n}))_{t\geq 0}$ is relatively compact for all $f\in\mathcal{A}$. Then, for given $f\in\mathcal{A}$, there exists a subsequence $(n_{k})_{k\in{\mathbb{N}}}$ such that $f(Y_{t}^{n_{k}})_{t\geq 0}$ converges to some limit $(X_{t}^{f})_{t\geq 0}$.
@@
Clearly, $(M_{f}^{\bar{n}_{k}})_{k}$, $(B_{f}^{\bar{n}_{k}})_{k}$, and $f(Y_{t}^{\bar{n}_{k}})_{k}$ are relatively compact for $f\in\mathcal{A}$. As in [18, Theorem 7.4.1], $(M_{f}^{n_{k}})_{k}$ is uniformly integrable for $f\in\mathcal{A}$. Consequently,